if sec 2 theta=4/3 then the general solution of theta

if sec2theta=4/3 then the general solution of theta

Grade:11

1 Answers

Aditya Gupta
4 years ago
we can rewrite it as
cos^2x= (sqrt(3)/2)^2= cos^2(pi/6)
we know that the general soln of cos^2(x)= cos^2(y) is x= npi ± y
so, gen soln is x= n*pi ± pi/6= pi*(n ± 1/6)= (6n ± 1)*pi/6
kindly approve :)))